OpenStudy (anonymous):

A open box with a volume of 1500 cm cube is to be constructed by taking a piece of cardboard 20 cm by 40 cm, cutting squares of side length x cm from each corner, and folding up the sides. Find the exact dimensions of the box.

OpenStudy (anonymous):

Solve the following volume expression, for x:\[(20-2x)(40-2x)x=1500\]There are three solutions for x: x=5 and x= (5/2)* (5 - Sqrt(13) )=3.48612 are the two of the three x solutions that are valid for this problem.